OpenStudy (anonymous):

(4t^2-21t+5)/(4t^2+15t-4) multiply (5t^2+19t-4)/(20t^2-9t+1)

OpenStudy (lgbasallote):

\[\frac{4t^2 - 121t + 5}{4t^2 + 15t - 4} \; \times \;\frac{5t^2 + 19t - 4}{20t^2 - 9t + 1}\] first step is to factor. i'll help you with the first step. \[\frac{(4t-1)(t - 5)}{(4t-1)(t+4)}\; \times \;\frac{(5t - 1)(t + 4)}{5t - 1)(4t - 1)}\] does that help?

OpenStudy (anonymous):

yeah so t-5/4t-1?

OpenStudy (lgbasallote):

i believe so